USING PROGRAM ERHAM TO ANALYZE HIGH-RESOLUTION INFRARED SPECTRA OF MOLECULES WITH INTERNAL ROTORS

\underline{P. GRONER}, Department of Chemistry, University of Missouri-Kansas City, Kansas City, MO 64110-2499; S. ALBERT, M. QUACK, Physical Chemistry, ETH Zürich, CH-8093 Zürich, Switzerland.

The effective rotational Hamiltonian for molecules with one or two periodic large-amplitude motions ^{a} implemented in program ERHAM has been adapted to enable prediction and least-squares fits of rotationally resolved lines in vibrationrotation spectra in the infrared region. The modified program is currently applied to assign the band of methyl formate at 925 cm^{−1} that has been measured at ETH in Zurich on the IFS125 Bruker prototype ZP 2001 FTIR spectrometer ^{b} at a resolution of 0.001 cm^{−1}. An external glass cell with an optical path length of 3 m contained the sample, and 150 interferograms were co-added. Right now it looks as if the splitting into A and E components were a little too small to be resolved sufficiently for positive identification.

aP. Groner, J. Chem. Phys. 107, 4483 (1997).

bS. Albert, M. Quack, ChemPhysChem 8, 1271 (2007)